7. (a) Consider the chlorination of methane, given below: 1 CH4(g) + 4 Cl2(g) 1 CCl4(g)...

7.

(a) Consider the chlorination of methane, given below:

1 CH4(g) + 4 Cl2(g) 1 CCl4(g) + 4 HCl(g)

If CH4(g) is decreasing at the rate of 0.490 mol/s, what are the rates of change of Cl2(g), CCl4(g), and HCl(g)?

Cl2(g)/t = ____ mol/s

CCl4(g)/t = _____mol/s

HCl(g)/t = _____mol/s

(b) The decomposition reaction given below:

2 IF5(g) 1 I2(g) + 5 F2(g)

is carried out in a closed reaction vessel. If the partial pressure of IF5(g) is decreasing at the rate of 256 torr/min, what is the rate of change of the total pressure in the vessel?

Ptot /t =____ torr/min

these answers are wrong:

Cl2(g)/t = 1.96 mol/s

CCl4(g)/t = -0.49 mol/s

HCl(g)/t = -1.96 mol/s
Answer #1

1 CH4(g) + 4 Cl2(g)-------> 1 CCl4(g) + 4 HCl(g)

-d[CH4]/dt   = -1/4 d[Cl2]/dt   = +d[CCl4]/dt   = +1/4 d[HCl]/dt

d[CH4]/dt   = -1/4d[Cl2]/dt

0.49          = -1/4d[Cl2]/dt

d[Cl2]/dt   = 1.96mole/sec

-d[CH4]/dt   = +d[CCl4]/dt

-0.49            =+d[CCl4]/dt

+d[CCl4]/dt   =- 0.49

-d[CH4]/dt   = +1/4 d[HCl]/dt

-0.49              = +1/4 d[HCl]/dt

d[HCl]/dt       = -1.96mole/sec